OpenStudy (anonymous):

find the values that make the expression undefined 3x-2 over y^2+49

OpenStudy (anonymous):

What makes this undefined?

OpenStudy (anonymous):

Well, its undefined when the denomintor is 0. Lets set the denominator to zero and solve for y

OpenStudy (anonymous):

y^2+49=0

OpenStudy (anonymous):

subtract 49 from both sides

OpenStudy (anonymous):

Then take the square root of both sides. leaving you with: y=?

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

but,. on second look, the denomintor never goes to zero, can you see why?

OpenStudy (anonymous):

because 7^2 = 49 +49 which then equals a positive number

OpenStudy (anonymous):

exactly, and thats why this function is never undefined

OpenStudy (anonymous):

can y= -(7^2)?

OpenStudy (anonymous):

Well, let me as you this, if i plug -7 into y^2, what do i get?

OpenStudy (anonymous):

ask*

OpenStudy (anonymous):

you get positive 49 right?

OpenStudy (anonymous):

so, because the y is squared, no matter what number you put in there, you will get a positive number

OpenStudy (anonymous):

thus, the denominator will never =0

OpenStudy (anonymous):

there is no value for which this expression is undefined

OpenStudy (anonymous):

but can't the (7^2) be multiplied first and then the negative multiplies it on the outside?

OpenStudy (anonymous):

no it is impossible ......there is no value for which this expression in undefined

OpenStudy (anonymous):

thats not possible, beacuse that would mean, that the negative sign would also have to affect the whole denominator: -(7^2+49)=-49-49=still defined